here is the question

Coffee and tea are substitutes. The demand for each depends on its own price as well as the price of its substitiute. Supply and demand curves are given below for each:-

Coffee Demand Qdc = 60-6Pc + 4Pt
Coffee Supply Qsc = 3Pc

Tea demanded Qdt = 20-2Pt + Pc
Tea Supply Qst =2Pt
Where Pc is the price of coffee and Pt is the price of Tea

Find the equilibrium prices and quantities for coffee and tea . Hint Both markets must be simultaneously in equilibrium.

Answers Are Pc =10 Qc =30 PT = 7.5 QT =15 ( but i don't know how they got to those answers ?)
1
Expert's answer
2018-12-28T05:15:11-0500

Coffee Demand Qdc = 60-6Pc + 4Pt

Coffee Supply Qsc = 3Pc

Tea demanded Qdt = 20-2Pt + Pc

Tea Supply Qst =2Pt

In equilibrium Qd = Qs.

60 - 6Pc + 4Pt = 3Pc

20 - 2Pt + Pc = 2Pt,

60 + 4Pt = 9Pc

+

20 - 4Pt = -Pc

=

8Pc = 80, Pc = 10, Qsc = 3*10 = 30.

20 - 4Pt = -10,

4Pt = 30, Pt = 7.5, Qst = 2*7.5 = 15.